OpenStudy (anonymous):

describe how yhe transformed function y = (x+3)^3 is obtaines from its parent function

OpenStudy (anonymous):

y=x^3 Parent function

OpenStudy (anonymous):

y=(x+a)^3 parent function shifted to the left a units y=(x-a)^3 parent function shifted to the right a units

OpenStudy (anonymous):

@doulikepiecauseidont thanks